Morgana Company identifies three activities in its manufacturing process: machine setups, machining, and inspections. Estimated annual overhead cost for each activity is $205,900, $411,200, and $58,500, respectively. The cost driver for each activity and the expected annual usage are number of setups 2,900, machine hours 25,700, and number of inspections 1,300.

Compute the overhead rate for each activity.

Machine setups $

  

per setup
Machining $

per machine hour
Inspections $

per inspection
Answer #1

Machine setups = $205,900 / 2,900 = $71 per setup

Machining = $411,200 / 25,700 = $16 per machine hour

Inspections = $58,500 / 1,300 = $45 per inspection

> accidentally downvoted but this is good thank you

Phuong Mon, Apr 19, 2021 2:02 AM